Identify each of the statements as being either TRUE or FALSE.
Young's modulus is defined as the ratio of stress to strain, consequently, it cannot be expressed in the same units as stress.

A: TrueB: False


The unit of pressure, the Pascal is dimensionally the same as newton/meter2.

A: TrueB: False
Answer #1

Stress \(=\frac{F}{A}\)

The units are \(\mathrm{N} / \mathrm{m}^{2}\)

Strain didn't have units.

Young's modulus \(=\frac{\text { Stress }}{\text {Strain}}\)

So the units of Young's modulus are nothing but the units of

Stress that is \(\mathrm{N} / \mathrm{m}^{2}\)


Hence the given statement is wrong, the answer is \(\overline{B . \text { False }}\)

The expression for the pressure is

\(P=\frac{F}{A}\)

The units are \(\mathrm{N} / \mathrm{m}^{2}\) and it is also called as Pascal.

So the Pascal is dimensionally the same as \(\mathrm{N} / \mathrm{m}^{2}\). Hence the given

statement is correct.

The answer is \(\overline{A \text { . True }}\)